### v2

# On the analytic continuation of Dirichlet series with missing digits

Abstract.
We study the Dirichlet series associated with the integers whose radix-$b$ representation misses certain (fixed) digits. The existence of a meromorphic continuation to the entire complex plane, which was already well-known as a general fact valid for $b$-automatic Dirichlet series, is proven anew from a representation as an everywhere defined series with good convergence properties. A generating function related to the residues on the real axis is shown to be the multiplicative inverse of the moment generating function for the associated Cantor set in the unit interval. This makes the (normalized) residues some sort of generalized Bernoulli numbers.
Key words and phrases:
Dirichlet series with missing digits, Riemann zeta function, meromorphic continuation, residues and poles, generating functions, Cantor sets, Bernoulli numbers
2020 Mathematics Subject Classification:
11M41 (Primary) 11A63, 11B85, 11M06, 11Y35, 33F05, 68R15 (Secondary)
1. Introduction
Throughout the paper, $\mathbb{N}$ is the set of non-negative integers and $\mathbb{Z}$ the set of relative integers. Let $b>1$ be an integer and $A\subset[\![0,b-1]\!]$ be a set of digits in base $b$, containing at least one non-zero digit. Let $N$ be the cardinality of $A$.
This paper is devoted to a study of some analytical properties of the “restricted” Dirichlet series:
$K(s)=\sideset{}{{}^{\prime}_{n>0}}{\sum}n^{-s}.$ (1)
Here, the symbol next to the summation sign means that only those integers are kept whose $b$-ary digits are all in $A$. We say that such integers are “admissible” (and the digits in $A$ are the “admissible digits”). The series should arguably be denoted $K_{b,A}(s)$ but we shorten the notation for convenience. If $N=b$, i.e. if every digit is admissible, then $K(s)$ is the defining series of the Riemann zeta function $\zeta(s)$.
This series is a member of the much wider class of $b$-automatic Dirichlet series. To explain this, rewrite $K(s)$ as $\sum_{n>0}c_{n}n^{-s}$ where $c_{n}=1$ if $n$ is admissible, $c_{n}=0$ otherwise (and we set $c_{0}=1$). The sequence $(c_{n})_{n\geq 0}$ is $b$-automatic (see [3, 2] for this notion; see also the recent paper [4] for a contemporary discussion of some aspects of Dirichlet series for $b$-automatic and also for the more general $b$-regular sequences). As such, thanks to a general theorem of Allouche, Mendès France and Peyrière ([2]; see also [9, Thm. 3.1] for $b$-regular sequences), $K(s)$ has a meromorphic extension to the entire complex plane, with candidate poles located in a finite union of left semi-lattices having $1$ and $2\pi i/\log b$ as periods.
The initial motivation for this paper came from another direction. In [7] we represented $K(s)$ in its half-plane of convergence $\Re s>\log_{b}N$ in terms of a geometrically convergent series (even in the case of the Riemann zeta function, such representation appears to be new), involving certain coefficients $(u_{m}(s))_{m\geq 0}$ which obey a finite recurrence which is stable numerically. They where first introduced for $s=1$ in [6], where a novel algorithm was developed for the numerical evaluation of $K(1)$ (the Kempner series [13, 5, 11, 6]).
In the first section of this paper, we prove that the series of [7] makes sense everywhere in the complex plane and that its terms (hence its remainders) are bounded by geometrically decreasing sequences, uniformly on compact sets (see Theorem 2.3 and Corollary 2.4 for the precise statement; a multiplicative factor is present to cancel the poles). This applies in particular to $\zeta(s)$, for which we had proven already the extension to $\Re s>0$ ([8]). A new technique is needed for the extension to the whole plane, and it is presented here in complete generality of an arbitrary set $A$ of admissible $b$-ary digits.
The case $N=1$ is very explicit and elementary and we discuss it briefly in the second section. We return next to the study with general $N$, and give in the third section a second proof of meromorphic continuation. We simply follow the same steps as in the far more general theorem of [2] regarding $b$-automatic Dirichlet series. Their method, and our adaptation–instantiation, is based upon an infinite functional equation expressing the value at $s$ in terms of the values at $s+1$, $s+2$, $s+3$, … .
Regarding the values at the non-positive integers, it will turn out that K(m)K(-m) is zero for all $m>0$ if 0 is an admissible digit. In the alternative, these values are less trivial and we compute their exponential generating function (this is done in the last section of this paper, Proposition 8.1).
Regarding the residues, we prove that, with $s_{0}=\log_{b}N$ the abscissa of convergence, there is a pole at $s_{m,k}=s_{0}-m+k2\pi i/\log(b)$, $m\in\mathbb{N}$, $k\in\mathbb{Z}$, if and only if both $s_{m,0}$ and $s_{0,k}$ are indeed poles: this follows from a proportionality of rows of (modified) residues, Proposition 7.3. We give a formula for the residue $\lambda_{0,k}$ at $s_{0,k}$ as a limit (Proposition 5.1), but do not study them further, preferring to focus on the residues $\lambda_{m,0}$ at the $s_{0}-m$ locations on the real axis. It proves convenient to replace them with some closely related quantities $\mu_{m,0}$ (see Equation (26)). We can define the latter quantities also in the case of the Riemann zeta function (for which the residues would be zero), and they turn out to be the Bernoulli numbers. Each choice of the integer $b>1$ gives its own linear recurrence characterizing them.
To conclude this summary, we mention a result which we find rather striking (from a “structural” point of view; we do not engage into any advanced analytics, though).
Consider the Cantor set of the numbers in $[0,1]$ whose (infinite) $b$-radix representation (or one of them, if there are two) uses only admissible digits (if $N=1$ it is reduced to a singleton and if $N=b$ it is the un-modified interval itself). It is well-known that there is a canonical probability measure $\mu$ on this set, singular to Lebesgue measure if $N<b$, whose moment generating function $\int_{[0,1]}\mathsf{e}^{tx}\mathrm{d}\mskip-1.0mu\mu(x)$ is the infinite product (which is an entire function):
$E(t)=\prod_{j=1}^{\infty}\frac{\sum_{a\in A}\mathsf{e}^{b^{-j}at}}{N}\;.$ (2)
In the case with $N=b$, a simple computation shows that the above infinite product, a priori depending on $b$, in fact indeed computes $(\mathsf{e}^{t}-1)/t=\int_{0}^{1}\mathsf{e}^{tx}\mathrm{d}x$.
Let us define $B(t)$ (for Bernoulli) as the multiplicative inverse of $E(t)$:
$B(t)=\prod_{j=1}^{\infty}\frac{N}{\sum_{a\in A}\mathsf{e}^{b^{-j}at}}\;.$ (3)
Theorem 7.4 states that $B(t)$ is the exponential generating function for the (modified) residues of $K(s)$ at the $s_{0}-m$ points, $m\geq 0$. In the case with $N=b$, this simply means that our ad hoc choice for the $\mu_{m,0}$ gives the sequence of Bernoulli numbers. So this makes the (normalized) residues in general a kind of generalization of Bernoulli numbers. Probably some authors have already envisioned (3) in this perspective, but we feel that the connection with the residues of the restricted Dirichlet series contributes to strengthen the analogy.
It would be interesting to check to what extent, some, or all, of our findings are examples of general facts for $b$-automatic Dirichlet series, or perhaps even more general ones. The literature on Dirichlet series, and the one on automatic sequences as well, are too large for the author to be able to provide here a curated list of the various works which could prove relevant to such endeavor. Thus, in addition to the papers already cited we mention here only a few, mostly related with the study of the sum-of-digits function or other summatory functions: [12, 9, 14, 10].
2. Representation by geometrically convergent series
We let $A_{1}=A\setminus\{0\}$, we already defined N=#AN=\#A, and we set $N_{1}=\#A_{1}\geq 1$. The length $\ell(n)$ of a natural integer $n\in\mathbb{N}$ is the number of digits of the minimal $b$-representation, so $\ell(0)=0$ (as the integer 0 is represented by the empty word) and $\ell(n)$ for $n>0$ is the smallest integer exponent $l$ such that $n<b^{l}$. Define for each positive integer $l$ the finite sum (“blocks”):
$K_{l}(s)=\sideset{}{{}^{\prime}_{\ell(n)=l}}{\sum}n^{-s}.$ (4)
We also define $K_{\leq\ell}(s)=\sum_{1\leq l\leq\ell}K_{l}(s)$, and $K_{\geq\ell}(s)=\sum_{l=\ell}^{\infty}K_{l}(s)$ (in case of convergence, at first), and similarly with other types of restrictions on the indices.
The cardinality of the set of admissible integers of length $l$ is $N_{1}N^{l-1}$. So, for $s=\sigma$ real, there holds $N_{1}N^{l-1}b^{-l\sigma}<K_{l}(\sigma)\leq N_{1}N^{l-1}b^{-(l-1)\sigma}$. Thus, the necessary and sufficient condition for the convergence of $K(\sigma)=\sum_{l=1}^{\infty}K_{l}(\sigma)$ is $b^{\sigma}>N$, i.e. $\sigma>s_{0}=\log_{b}N$.
A general basic result (Jensen lemma, based upon a partial summation, see [17, Chap. 4]), which underlies the notion of abscissa of convergence, says that as a consequence, the Dirichlet series (1) defining $K(s)$ diverges whenever $\Re(s)<s_{0}$. The abscissa of convergence is thus $s_{0}$ ([15]; see also [16] for more generality, and [4] for a still wider perspective). Let us mention this complement:
Proposition 2.1.
For any $s=s_{0}+it$, $t\in\mathbb{R}$, the Dirichlet series $K(s)$ diverges.
Proof.
Suppose to the contrary that it converges for some $s=s_{0}+it$. Let $n_{1}<n_{2}<\dots<n_{q}<\dots$ be the positive admissible integers. Let $r_{p}=\sum_{q=p}^{\infty}n_{q}^{-s}=o(1)$. We do a summation by parts:
$S_{p}\coloneq\sum_{q=1}^{p}n_{q}^{-s_{0}}=\sum_{q=1}^{p}(r_{q}-r_{q+1})n_{q}^{it}=r_{1}n_{1}^{it}+\sum_{q=2}^{p}r_{q}\bigl(n_{q}^{it}-n_{q-1}^{it}\bigr)-r_{p+1}n_{p}^{it}.$
From $n_{q}^{it}-n_{q-1}^{it}=\int_{n_{q-1}}^{n_{q}}itx^{it-1}\mathrm{d}x$, or from $|\mathsf{e}^{i\theta}-1|\leq|\theta|$, we get $\big|n_{q}^{it}-n_{q-1}^{it}\bigr|\leq|t|\log\frac{n_{q}}{n_{q-1}}$. The summation on the right-hand side, restricted to the indices $q$ such that $p_{0}<q\leq p$ with some fixed $p_{0}$ is thus bounded by $|t|(\sup_{q>p_{0}}|r_{q}|)\log(n_{p}n_{p_{0}}^{-1})$. Bounding trivially the finitely many other terms, it follows:
$\lim_{p\to\infty}\frac{S_{p}}{\log n_{p}}=0.$
But this is wrong. To see it, take $\ell\geq 1$ and $p_{\ell}=N_{1}+N_{1}N+\dots+N_{1}N^{\ell-1}$ so that $n_{p_{\ell}}$ is the largest admissible integer which is $b^{\ell}$. Thus $n_{p_{l}}<b^{\ell}$. And $S_{p_{\ell}}>N_{1}b^{-s_{0}}+N_{1}Nb^{-2s_{0}}+N_{1}N^{2}b^{-3s_{0}}+\dots+N_{1}N^{\ell-1}b^{-\ell s_{0}}=\ell N_{1}N^{-1}$. Hence, $S_{p_{\ell}}$ is not $o(\log n_{p_{\ell}})$. ∎
In [7], the following quantities are defined for $\Re s>s_{0}$, and $m\geq 0$:
$u_{m}(s)=0^{m}+\sum_{l=1}^{\infty}\sum_{(a_{1},\dots,a_{l})\in A^{l}}\bigl(a_{1}b^{-1}+\dots+a_{l}b^{-l}\bigr)^{m}\,b^{-ls}.$ (5)
They are the moments of a certain complex measure on the set of $b$-words, but we will not use this vocabulary here. The defining series for $u_{m}(s)$ converges if and only if it converges absolutely if and only if $\Re s>s_{0}$ (use that the sum of powers is bounded below by $N_{1}N^{l-1}b^{-m}$ and above by $N^{l}$). One has
$u_{0}(s)=\frac{b^{s}}{b^{s}-N}\,$,
and the recurrence relation ([7, Prop. 1]):
$m\geq 1\implies(b^{s+m}-N)u_{m}(s)=\sum_{j=1}^{m}\binom{m}{j}(\sum_{a\in A}a^{j})u_{m-j}(s),$ (6)
which will be the basis for our further investigations in this section.
We use the following notation for power sums over the admissible digits:
$\gamma_{j}=\sum_{a\in A}a^{j}.$
In particular, $\gamma_{0}=N$. We proved ([7, Thm. 1]) the following representation of $K(s)$ in its half-plane of convergence. For any choice of a level $\ell\geq 1$, and any $s$ with $\Re s>s_{0}$:
$K(s)=\sideset{}{{}^{\prime}_{0<n<b^{\ell-1}}}{\sum}\frac{1}{n^{s}}+\sum_{m=0}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}u_{m}(s)\sideset{}{{}^{\prime}_{b^{\ell-1}\leq n<b^{\ell}}}{\sum}\frac{1}{n^{m+s}}.$ (7)
We will simply use $\ell=2$ in the sequel (geometric convergence may not apply if $\ell=1$, see [7, Thm. 1] for details). The notation $(s)_{m}$ is the Pochhammer symbol (partial ascending factorial with $m$ terms).
From (6) and $u_{0}(s)=b^{s}/(b^{s}-N)$, it is clear that all the $u_{m}(s)$ extend meromorphically to the entire complex plane, with simple candidate poles at the points of the left semi-lattice:
$\Delta=\{s_{m,k}=s_{0}-m+k\frac{2\pi i}{\log b},m\in\mathbb{N},k\in\mathbb{Z}\}.$
In the case with $N=b$ (and also $N=1$, see the next section) there are explicit alternative formulas for the $u_{m}(s)$’s. If $N=b$ one sees that they have poles only for $\Re s\in\{1,0,-1,-3,-5,\dots\}$. The identical poles coming from multiple $u_{m}(s)$’s conspire to cancel out, and in the end only the pole at $s=1$ remains for $\zeta(s)$ itself…
Our aim in this section is to prove that the series at the right-hand side of (7) converges everywhere in $\mathbb{C}\setminus\Delta$, thus providing the analytic continuation of $K(s)$. We will prove this in a slightly stronger form, so that it will be shown that $K(s)$ is meromorphic with only simple poles. The proof is elementary but needs some notation and care. The key ingredient is the following Lemma:
Lemma 2.2.
Let $p$ be a positive integer. Let $0<\eta$. Let $(c_{m})_{m\geq 0}$ be the sequence of positive real numbers, depending upon $p$ and $\eta$, and defined by the conditions:
$c_{0}=c_{1}=\dots=c_{p-1}=1$
$m\geq p\implies(b^{1+\eta+m-p}-N)c_{m}=\sum_{j=1}^{m}\binom{m}{j}\gamma_{j}c_{m-j}.$
Let $f=\max A$ and $\lambda=f/(b-1)$. There exists an (explicit) continuous function $C_{p}$ on the positive real numbers such that $c_{m}\leq C_{p}(\eta)\lambda^{m}m^{p-1-\eta}$ for every $m\geq p$.
Remark 2.1.
For $N<b$ we could take $\eta=0$, as the only problem with $\eta=0$ is with the definition of $c_{p}$ if bN=0b-N=0.
Proof of Lemma 2.2.
We let $\tau=1+\eta$. Let $n\geq 0$ and $m=n+p$. We split the binomial sum on the right-hand side of the recurrence as $A_{n}+B_{n}$ where $B_{n}$ stands for the contributions of the last $p$ terms. We will take care of $B_{n}$ later. The recurrence relation is:
$(b^{\tau+n}-N)c_{p+n}=\sum_{j=1}^{n}\binom{p+n}{j}\gamma_{j}c_{p+n-j}+B_{n}.$
Let us now make the definition, for $n\geq 0$:
$d_{n}=\frac{(\tau+1)_{n}}{(p+n)!}c_{p+n}=\frac{(\tau+1)\dots(\tau+n)}{(p+n)!}c_{p+n}.$
We compute:
$(b^{\tau+n}-N)d_{n}$ $=\frac{(\tau+1)_{n}}{(p+n)!}A_{n}+\frac{(\tau+1)_{n}}{(p+n)!}B_{n}$
$=\sum_{j=1}^{n}\frac{(p+n)!}{j!(p+n-j)!}\gamma_{j}\frac{(\tau+1)_{n}}{(p+n)!}\frac{(p+n-j)!}{(\tau+1)_{n-j}}d_{n-j}+\frac{(\tau+1)_{n}}{(p+n)!}B_{n}$
$=\sum_{j=1}^{n}\frac{(\tau+n)\dots(\tau+n-j+1)}{j!}\gamma_{j}d_{n-j}+\frac{(\tau+1)_{n}}{(p+n)!}B_{n}.$
We claim that
$\sum_{j=1}^{n}\frac{(\tau+n)\dots(\tau+n-j+1)}{j!}\gamma_{j}\lambda^{-j}\leq b^{\tau+n}-N.$ (8)
The argument (for which $\tau\geq 1$ is important) was given in [7, Proof of Prop. 5]. For the convenience of the reader, we repeat it partially here. First observe that $\gamma_{j}$ can be written as $\sum_{a\in A_{1}}a^{j}$ as $j\geq 1$. We have to evaluate:
$\begin{split}\sum_{j=1}^{n}\frac{(\tau+n)\dots(\tau+n-j+1)}{j!}\sum_{a\in A_{1}}\bigl(\frac{(b-1)a}{f}\bigr)^{j}\\ \leq\sum_{j=1}^{n}\frac{(\tau+n)\dots(\tau+n-j+1)}{j!}\sum_{a=f,f-1,\dots,f-N_{1}+1}\bigl(\frac{(b-1)a}{f}\bigr)^{j}.\end{split}$
The following calculus lemma ([7, Lem. 3]) holds: Let $\sigma\geq 1$ be a real number, $n$ a non-negative integer, and $x$ a non-negative real number. Then
$\sum_{j=1}^{n}\frac{(n+\sigma)\dots(n+\sigma-j+1)}{j!}x^{j}\leq(1+x)^{n+\sigma}-1-x^{n+\sigma}.$
Thus, if b1=fb-1=f we obtain directly a telescopic sum as upper bound. If b1>fb-1>f, one needs only make a simple additional remark to again reduce to a telescopic sum, see [7, Proof of Prop. 5]. This gives us:
$\textit{currently evaluated sum}\leq b^{\tau+n}-N_{1}-1\leq b^{\tau+n}-N,$
which validates claim (8). Let now, for $k\geq 0$, $D_{k}=\max_{0\leq j\leq k}\lambda^{-j}d_{j}$ (in particular, $D_{0}=d_{0}=\frac{c_{p}}{p!}$), so that $d_{n-j}\leq\lambda^{n}\lambda^{-j}D_{n-1}$ for $1\leq j\leq n$. We get, using (8):
$(b^{n+\tau}-N)d_{n}$ $\leq\lambda^{n}(b^{\tau+n}-N)D_{n-1}+\frac{(\tau+1)_{n}}{(p+n)!}B_{n}$
$\lambda^{-n}d_{n}$ $\leq D_{n-1}+\frac{(\tau+1)_{n}}{(p+n)!}\frac{\lambda^{-n}B_{n}}{b^{n+\tau}-N}$
$D_{n}$ $\leq D_{n-1}+\frac{(\tau+1)_{n}}{(p+n)!}\frac{\lambda^{-n}B_{n}}{b^{n+\tau}-N}\leq\sum_{n=0}^{\infty}\frac{(\tau+1)_{n}}{(p+n)!}\frac{\lambda^{-n}B_{n}}{b^{n+\tau}-N}.$
(The $n=0$ term is indeed $D_{0}=d_{0}$). Let $D^{p}(\tau)$ be the series on the right-hand side. We show that it is a continuous function of $\tau>0$. There holds (replacing in the last step (n+pk)!(n+p-k)! by $(n+1)!$):
$B_{n}=\sum_{j=n+1}^{n+p}\binom{n+p}{j}\gamma_{j}=\sum_{k=0}^{p-1}\binom{n+p}{k}\gamma_{n+p-k}\leq N_{1}\sum_{k=0}^{p-1}\frac{(n+p)!}{k!(n+1)!}f^{n+p-k},$
[…]
$\sum_{n=0}^{\infty}\frac{(\tau+1)_{n}}{(n+1)!}\frac{(b-1)^{n}}{b^{n+\tau}-N}.$
Suppose $1<\tau_{0}\leq\tau\leq\tau_{1}<\infty$. Then the above is bounded termwise by
$\sum_{n=0}^{\infty}\frac{(\tau_{1}+1)_{n}}{(n+1)!}\frac{(b-1)^{n}}{b^{n+\tau_{0}}-N}.$
The latter series converges by the ratio test. Consequently $D^{p}(\tau)<\infty$ with locally uniform convergence, hence it is a continuous function such that $D_{n}\leq D^{p}(\tau)$ for every $n\geq 0$. Hence, $d_{n}\leq D^{p}(\tau)\lambda^{n}$ and consequently
$m\geq p\implies c_{m}$ $\leq D^{p}(\tau)\frac{m!}{(\tau+1)_{m-p}}\lambda^{m-p}$
$\implies c_{m}(m-p)^{\tau}m^{-p}\lambda^{-m}$ $\leq\lambda^{-p}D^{p}(\tau)\frac{(m-p)!(m-p)^{\tau}}{(\tau+1)_{m-p}}$
The last equation is useful only for $m\geq p+1$. Let now $\alpha_{n}=n!n^{\tau}/(\tau+1)_{n}$ for $n\geq 1$. As is well-known, $\alpha_{n}$ converges toward $\Gamma(\tau+1)$, and of course this is uniform for $\tau$ bounded. In fact, the sequence $(\alpha_{n})$ is increasing, and thus bounded above by $\Gamma(\tau+1)$. To check that, i.e. to prove that
$\frac{(n+1)(n+1)^{\tau}n^{-\tau}}{n+1+\tau}\geq 1,$
[…]
$\log(n+1)-\log(n)-\frac{1}{n+1+\tau}=\int_{n}^{n+1}\frac{\mathrm{d}t}{t}-\frac{1}{n+1+\tau}>0.$
So, for $m>p$ we obtain
$c_{m}\leq\bigl(1-\frac{p}{m}\bigr)^{-\tau}\lambda^{-p}D^{p}(\tau)\Gamma(\tau+1)\lambda^{m}m^{p-\tau}\leq\bigl(p+1\bigr)^{\tau}\lambda^{-p}D^{p}(\tau)\Gamma(\tau+1)\lambda^{m}m^{p-\tau}$
And as $c_{p}\leq D^{p}(\tau)p!$, the above upper bound holds also for $m=p$. As $\tau=1+\eta$, this completes the (somewhat long) proof of Lemma 2.2. ∎
We can now establish straightforwardly our main objective. Recall $\lambda=f(b-1)^{-1}$, $f=\max A$.
Theorem 2.3.
Let $\alpha(s)$ be the entire function $\prod_{n=0}^{\infty}(1-b^{-s-n}N)$. One has, uniformly on every compact subset of the complex plane:
$\alpha(s)\frac{(s)_{m}}{m!}u_{m}(s)=_{m\to\infty}O\bigl(\lambda^{m}m^{-1}\bigr).$
Proof.
Let $Q\subset\mathbb{C}$ be some non-empty compact subset and choose $p$ a positive integer such that $\Re s\geq 2-p$ for every $s\in Q$. For any $s\in Q$, we let $\eta(s)=\Re s+p-1\geq 1$. The analytic coefficients $v_{m}(s)=\alpha(s)u_{m}(s)$, $m\geq 0$, verify the recurrence
$(b^{m+s}-N)v_{m}(s)=\sum_{j=1}^{m}\binom{m}{j}\gamma_{j}v_{m-j}(s).$
The sequence of positive real numbers $(c_{m})$ from Lemma 2.2, for this choice of $p>0$ and $\eta=\eta(s)\geq 1$, verify the recurrence, starting at $m=p$,
$m\geq p\implies(b^{m+\Re s}-N)c_{m}=\sum_{j=1}^{m}\binom{m}{j}\gamma_{j}c_{m-j}\,$,
and with initial conditions $c_{0}=\dots=c_{p-1}=1$.
Observe that $|b^{m+s}-N|\geq b^{m+\Re s}-N\geq b^{2}-N>0$ for any $m\geq p$. Thus, as the binomial coefficients and power sums are positive, an induction proves that
$\forall m\geq 0\quad|v_{m}(s)|\leq c_{m}\max_{0\leq j\leq p-1}\|v_{j}(s)\|_{\infty},$
where the sup-norms are computed on $Q$. Hence, according to the conclusion of Lemma 2.2, and setting
$M_{Q}=\max_{0\leq j\leq p-1}\|\alpha(s)u_{j}(s)\|_{\infty},\qquad M_{Q}^{\prime}=M_{Q}\sup_{s\in Q}C_{p}(\eta(s)),$
we get, using $p-1-\eta(s)=-\Re s$:
$\forall m\geq p\quad|\alpha(s)u_{m}(s)|\leq M_{Q}C_{p}(\eta(s))\lambda^{m}m^{p-1-\eta(s)}\leq M_{Q}^{\prime}\lambda^{m}m^{-\Re s}.$
It is known that
$\frac{m^{1-s}(s)_{m}}{m!}\mathop{\longrightarrow}\limits_{m\to\infty}\frac{1}{\Gamma(s)},$
uniformly on every compact subset of the complex plane. This completes the proof of the uniform $O(\lambda^{m}m^{-1})$ estimate for $\alpha(s)(s)_{m}(m!)^{-1}u_{m}(s)$. ∎
Corollary 2.4.
The series, whose terms are entire functions,
$\sum_{m=0}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}\alpha(s)u_{m}(s)\sum_{\begin{subarray}{c}a_{1}\in A_{1}\\[1.5pt] a_{2}\in A\end{subarray}}(a_{1}b+a_{2})^{-s-m},$ (9)
converges everywhere in the complex plane and defines an entire function which gives the analytic continuation of $\alpha(s)\Bigl(K(s)-\sum_{a\in A_{1}}a^{-s}\Bigr)$. Consequently, $K(s)$ is a meromorphic function in the complex plane with candidate simple poles at the zeros of $\alpha(s)=\prod_{n=0}^{\infty}(1-b^{-s-n}N)$.
Proof.
Immediate consequence of Theorem 2.3: indeed, equation (15) tells us that for $\Re s>s_{0}$ the series of Equation (9) converges toward $\alpha(s)\Bigl(K(s)-\sum_{a\in A_{1}}a^{-s}\Bigr)$. As it converges locally uniformly everywhere it provides the analytic continuation. ∎
3. The case with $N=1$
Suppose $N=1$. There is only one admissible integer of length $l$, and it is $\lambda(b^{l}-1)$, with $\lambda=f(b-1)^{-1}$ where $f>0$ is the sole admissible digit. So,
$K(s)=\Bigl(\frac{b-1}{f}\Bigr)^{s}\sum_{l=1}^{\infty}\frac{1}{(b^{l}-1)^{s}}.$
We will assume from here on in this section that f=b1f=b-1, as the general case reduces to this one. There is a well-known identity valid for $\Re s>0$ (but we could not locate a standard reference):
$\sum_{l=1}^{\infty}\frac{1}{(b^{l}-1)^{s}}=\sum_{j=0}^{\infty}\frac{(s)_{j}/j!}{b^{s+j}-1}.$ (10)
To prove this, one only needs by analyticity, to do it for $s$ real and positive. It then follows from interverting summations in a double sum with positive terms, obtained from using the binomial series. The right-hand side in (10) converges everywhere in $\mathbb{C}\setminus\Delta$, where $\Delta$ is the left semi-lattice $\{-m+k2\pi i(\log b)^{-1},m\in\mathbb{N},k\in\mathbb{Z}\}$ (and at a candidate pole, removing the term where it intervenes, we have an analytic function).
Consequently, all the $s_{m,k}=-m+k2\pi i(\log b)^{-1},m\in\mathbb{N},k\in\mathbb{Z}$, are simple poles and the residues are:
$\lambda_{m,k}=\frac{(-m+k\frac{2\pi i}{\log b})(-m+1+k\frac{2\pi i}{\log b})\dots(-1+k\frac{2\pi i}{\log b})}{m!\,\log b},$ (11)
For $m=0$ the numerator is to be interpreted as being $1$. For $k=0$ one has $\lambda_{m,0}=(-1)^{m}(\log b)^{-1}$.
One may wonder if the series (10) is the one given by our Corollary 2.4, which states the validity everywhere of Equation (7). The answer is no! To start with, Equation (7) has a parameter $\ell$, which can be any positive integer, and it does a resummation of the original Dirichlet series from the left-hand side of (10) only starting at $l=\ell$.
We need to explain what the coefficients $(u_{m}(s))_{m\geq 0}$ are for this case. Let us denote them $w_{m}(s)$. From Equation (6), they verify the recurrence
$(b^{s+m}-1)w_{m}(s)=\sum_{j=1}^{m}\binom{m}{j}(b-1)^{j}w_{m-j}(s),$
with $w_{0}(s)=b^{s}(b^{s}-1)^{-1}$. These coefficients are known. Actually we only have to use formula (5) which gives the original definition of the sequence $(u_{m}(s))_{m\geq 0}$. For $\Re s>0$:
$w_{m}(s)=0^{m}+\sum_{l=1}^{\infty}(1-b^{-l})^{m}b^{-ls}=\sum_{l=0}^{\infty}(1-b^{-l})^{m}b^{-ls}$
Thus, we have the explicit finite Lambert-type sums here:
$w_{m}(s)=\sum_{j=0}^{m}(-1)^{j}\binom{m}{j}\frac{b^{s+j}}{b^{s+j}-1}=0^{m}+\sum_{j=0}^{m}(-1)^{j}\binom{m}{j}\frac{1}{b^{s+j}-1}.$ (12)
We can then enjoy exercises such as checking, for every $\ell\geq 1$:
$\sum_{1\leq l<\ell}\frac{1}{(b^{l}-1)^{s}}+\sum_{m=0}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}\frac{w_{m}(s)}{(b^{\ell}-1)^{m+s}}=\sum_{j=0}^{\infty}\frac{(s)_{j}/j!}{b^{s+j}-1},$
which is what Equation (7) and Corollary 2.4 say, combined with (10).
4. Analytic continuation (second method)
In this section we prove again that $K(s)=\sum^{\prime}_{n>0}n^{-s}=\sum_{l=1}^{\infty}K_{l}(s)$ admits a continuation to the whole complex plane as a meromorphic function. The method is the general one from [2] (see also the earlier paper [1]). Our account will be very detailed, probably too much!
Two cases often require special provisions: $N=1$ and $N=b$. They correspond to very explicit Dirichlet series:
  • If $N=b$, there are no restrictions and $K(s)=\zeta(s)$, the Riemann zeta function,
  • If $N=1$ then $K(s)=((b-1)/f)^{s}\sum_{l=1}^{\infty}(b^{l}-1)^{-s}$ where $f>0$ is the sole admissible digit. This was the topic of the previous section.
The first pole on the real axis is located for them at an integer point $s_{0}$: $s_{0}=1$ for $\zeta(s)$ and $s_{0}=0$ for $N=1$. All other cases are with $0<s_{0}<1$.
As in [2] (where the analogous quantity is denoted $F(s)$, whereas our $K(s)$ is analogous to their $G(s)$), it proves convenient to also consider the variant
$H(s)=\sideset{}{{}^{\prime}_{n>0}}{\sum}(n+1)^{-s}.$
The abscissa of convergence is again $\log_{b}N$, the counting argument given near the start of the first section applying almost identically. In the case with $A=\{b-1\}$, we simply have $H(s)=\frac{1}{b^{s}-1}$, which has poles only on the line $\Re s=0$. But, as we have seen, $K(s)$ in this case has poles on all lines $\Re s=-m$, $m\in\mathbb{N}$.
We establish that K(s)H(s)K(s)-H(s) is analytic for $\Re s>s_{0}-1$. Consider the simple estimate (assuming $\sigma=\Re s\geq-1$):
$\bigl|n^{-s}-(n+1)^{-s}\bigr|=\left|\int_{n}^{n+1}\frac{s\,\mathrm{d}t}{t^{s+1}}\right|\leq|s|n^{-\sigma-1}.$
This bound implies that $\sum^{\prime}\bigl(n^{-s}-(n+1)^{-s}\bigr)$ converges to an analytic function in the half-plane $\Re s>s_{0}-1$. So $K(s)$ and $H(s)$ have the same poles and residues in this half-plane.
We observe that admissible integers of length $l+1$ are of the shape $bn+a$, with $n$ admissible of length $l$ and $a\in A$. We now replace $(bn+a)^{-s}$ in $K_{l+1}(s)$ by its binomial series expansion in powers of $a/(bn)$. This leads to a representation of $K_{l+1}(s)$ in terms of the $K_{l}(s+m)$’s, $m\geq 0$. This step is also the starting point in the Baillie algorithm [5].
$(bn+a)^{-s}$ $=\sum_{m=0}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}(bn)^{-s-m}a^{m},$
$K_{l+1}(s)$ $=\sum_{m=0}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}b^{-s-m}\gamma_{m}K_{l}(s+m),$ (13)
[…]
The Pochhammer symbol $(s)_{m}$ is defined to be $1$ for $m=0$ (this rule applies to all empty products). Except for $m=0$, $\gamma_{m}$ is also $\sum_{a\in A_{1}}a^{m}$.
We now show that the series (13) is absolutely convergent, uniformly for $s=\sigma+it$ in any compact subset of $\{\Re s>\log_{b}N\}$. First we take note of the bound
$\frac{|(s)_{m}|}{(\sigma)_{m}}\leq\frac{\Gamma(\sigma)}{|\Gamma(s)|}.$
Indeed, $\Gamma(s)(s)_{m}=\Gamma(s+m)$ and we can use the integral representation to get $|\Gamma(s+m)|\leq\Gamma(\sigma+m)$ if $\sigma+m>0$. Here, $\sigma>0$ and $(\sigma)_{m}>0$. Thus, the series in (13) is bounded termwise by
$\frac{\Gamma(\sigma)}{|\Gamma(s)|}\sum_{m}\frac{(\sigma)_{m}}{m!}b^{-\sigma-m}\gamma_{m}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}n^{-\sigma-m}=\frac{\Gamma(\sigma)}{|\Gamma(s)|}\mathop{\sum\nolimits^{\prime}}\limits_{\begin{subarray}{c}\ell(n)=l\\ a\in A\end{subarray}}(bn-a)^{-\sigma}.$
Call $\alpha_{l}$, say, the finite sum at the right-hand side. As $bn-a\geq b^{l-1}$ if $n\geq b^{l-1}$, we have certainly $\alpha_{l}\leq\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}Nb^{-(l-1)\sigma}=N_{1}\cdot N^{l-1}\cdot N\cdot b^{-(l-1)\sigma}$. Hence the series $\sum_{l\geq l}\alpha_{l}$ converges, as soon as $\sigma>\log_{b}N$.
Hence, for $\Re s=\sigma>\log_{b}N$, we can sum the identity (13) from $l=1$ to infinity and intervert the summations over $l$ and over $m$, obtaining:
$K_{>\ell}(s)=Nb^{-s}K_{\geq\ell}(s)+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}b^{-s-m}\gamma_{m}K_{\geq\ell}(s+m).$
Adding the block $K_{\ell}(s)$ to both sides, and moving $K_{\geq\ell}(s)$ to the left-hand side, we obtain:
$\bigl(1-Nb^{-s}\bigr)K_{\geq\ell}(s)=K_{\ell}(s)+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}b^{-s-m}\gamma_{m}K_{\geq\ell}(s+m).$ (14)
Equation (14) is for now established only for $\Re s=\sigma>\log_{b}N$.
From here on, let $s_{0}=\log_{b}N$.
To prepare for the upcoming inductive reasoning, we observe that for any compact set $Q$ in the complex plane, if we choose $m_{0}$ such that $m_{0}+Q\subset\{z,\Re z>s_{0}\}$, the terms of the above series starting with $m=m_{0}$ are defined (as evaluations are done in the half-plane $\Re s>s_{0}$) and this sub-series converges absolutely and uniformly. Indeed, we can bound $K_{\geq\ell}(s+m)$ by $K(m_{0}+\inf(\{\Re z,z\in Q\}))$, and $\gamma_{m}$ by $N(b-1)^{m}$, and then apply the (uniform) ratio test.
In the next proposition, the cases $N=b$ (i.e. when $K(s)=\zeta(s)$) and $N=1$ both have their own specificities, but we do not handle them here separately.
Proposition 4.1.
Let $b>1$ be an integer and $A$ be some subset of [[0,b1]][\![0,b-1]\!], not empty and not reduced to $\{0\}$, of cardinality $N$. The associated restricted Dirichlet series $K(s)$ admits a meromorphic continuation, still denoted $K(s)$, to the entire complex plane. It verifies the identity
$\bigl(1-Nb^{-s}\bigr)K(s)=\sum_{a\in A_{1}}a^{-s}+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}b^{-s-m}\gamma_{m}K(s+m),$ (15)
where $(s)_{m}$ is the Pochhammer symbol and the power sums $\gamma_{m}$ are defined as:
$\gamma_{m}=\sum_{a\in A}a^{m}.$ (16)
The function
$\Lambda_{K}(s)=\prod_{m=0}^{\infty}(1-Nb^{-s-m})K(s),$ (17)
is entire. The function $K(s)$ has a simple pole at $s_{0}=\log_{b}N$, with a positive residue $\lambda_{0}$. It has only simple poles in the complex plane, which are among the $s_{m,k}=s_{0}-m+k\frac{2\pi i}{\log b}$, $m\in\mathbb{N}$, $k\in\mathbb{Z}$.
Remark 4.1.
As we already mentioned this is only a very special instance of the general result from [2].
Proof.
Let us first suppose that $\Re s>s_{0}-1$. So the series in (15) (which is the special case $\ell=1$ for Equation (14), whose validity is known for $\Re s>s_{0}$) converges and defines an analytic function in that half-plane. Hence, $K(s)$ admits a meromorphic continution to it, with potential simple poles at the zeros $s_{0,k}$ of the equation $b^{s}=N$, which all have their real part equal to $s_{0}$. As $\lim_{\sigma\to s_{0}^{*}}K(\sigma)=+\infty$, there really is a pole at $s_{0}$ and its residue $\lambda_{0,0}$ is positive.
Suppose by induction on $m\geq 1$ that we have proven analytic continuation as a meromorphic function with simple poles to $\Re s>s_{0}-m$. This is known for $m=1$. The right-hand side of Equation (15) defines a meromorphic function in the half-plane $\{\Re s>s_{0}-m-1\}$. In the extended strip $1+s_{0}-m<\Re s<s_{0}$, the division by $(1-Nb^{-s})$ brings no extra singularity. So this equation defines the analytic continuation, with simple poles located at the roots of the equations $b^{s}=N$, $b^{s+1}=N$, …, $b^{s+m}=N$. By induction, this is true for every $m\geq 1$.
The infinite product $\prod_{m=0}^{\infty}(1-Nb^{-s-m})$ is absolutely and uniformly convergent on any compact set, after having removed finitely many factors so that its terms have no zeros there, so it defines an entire function with simple zeros at the $s_{m,k}$, $m\in\mathbb{N}$, $k\in\mathbb{Z}$. Hence $\Lambda_{K}(s)$ from (17) is entire. And we have proven that (15) holds everywhere in the complex plane. ∎
Remark 4.2.
In the case of the Riemann zeta function, $N=b$, the statement places potential poles at $1-\mathbb{N}+\frac{2\pi i}{\log b}\mathbb{Z}$. Using $b=2$ and $b=3$ and the linear independence over the rationals of $(\log 2,\log 3)$, we can reduce the potential pole locations to 11-$\mathbb{N}$. We shall provide another proof later, and only then address the urgent task of showing that $1$ is the only pole!
Prior to engaging into a more detailed discussion of the poles and residues, we state here the analytic continuation for $H(s)=\sum^{\prime}_{n>0}(n+1)^{-s}$.
Proposition 4.2.
The modified restricted Dirichlet series $H(s)=\sum^{\prime}_{n>0}(n+1)^{-s}$ admits a meromorphic continuation, still denoted $H(s)$, to the entire complex plane. The function
$\Lambda_{H}(s)=\prod_{m=0}^{\infty}(1-Nb^{-s-m})H(s).$ (18)
is entire. The relation
$\bigl(1-Nb^{-s}\bigr)H(s)=\sum_{a\in A_{1}}(a+1)^{-s}+\sum_{m=1}^{\infty}\frac{(s)_{m}}{m!}b^{-s-m}\gamma_{m}^{\prime}H(s+m)$ (19)
holds at all $s$ such that $s\neq s_{m,k}=\log_{b}N-m+k\frac{2\pi i}{\log b}$, $m\geq 0$, $k\in\mathbb{Z}$. Here the modified power sums $\gamma_{m}^{\prime}$ are defined as
$\gamma_{m}^{\prime}=\sum_{a\in A}(b-1-a)^{m}.$ (20)
Proof.
Let $l\geq 1$. Let $n_{1}$ be an admissible integer of length $l+1$, $n_{1}=bn+a$, $a\in A$. We use the decompostion
$n_{1}+1=bn+a+1=b(n+1)-(b-1-a),$
and binomial series to compute $(n_{1}+1)^{-s}$. Summing over the admissible digits $a\in A$ will be expressed via the variant power sums $\gamma_{m}^{\prime}$. The power series is the one of $(1-t)^{-s}$ which absorbs the $(-1)^{m}$ arising in $(1+t)^{-s}$. The reasoning leading to (19) is then almost identical to the one which gave (15). Details left to the reader. ∎
All contents of this section are but an illustration of long since known facts [2].
5. Residues (I)
Let us abridge the residue $\lambda_{0,0}$ at $s_{0}$ to $\lambda_{0}$. As a consequence of (15), we obtain
$\lambda_{0}\log b=\sum_{a\in A_{1}}a^{-s_{0}}+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s_{0})_{m}}{m!}N^{-1}b^{-m}\gamma_{m}K(s_{0}+m).$ (21)
The above equation does not make it particularly easy to check the positivity, but it is possible. If $s_{0}=0$ (i.e. $N=1$), the series above vanishes identically, the set $A_{1}$ is a singleton $\{a\}$ and $\lambda_{0}\log b=1$. We suppose $s_{0}>0$ from now on for this discussion. Let us first observe that (as is well-known) the individual series:
$(bn+a)^{-s_{0}}-(bn)^{-s_{0}}=\sum_{m=1}^{\infty}(-1)^{m}\frac{(s_{0})_{m}}{m!}\frac{a^{m}}{(bn)^{s_{0}+m}}$
have the property that the remainders have the same sign as their first term (as $s_{0}>0$). This property is inherited by the series from (21). So, as $s_{0}>0$, there holds
$\lambda_{0}\log b>\sum_{a\in A_{1}}a^{-s_{0}}-\frac{s_{0}\gamma_{1}}{Nb}K(s_{0}+1).$
The same principle applied at $s=s_{0}+1$ in (15), with $\ell=1$ gives
$(1-b^{-1})K(s_{0}+1)<\sum_{a\in A_{1}}a^{-s_{0}-1}.$
Combining the two estimates we get
$\lambda_{0}\log b>\sum_{a\in A_{1}}a^{-s_{0}}-\frac{s_{0}\gamma_{1}}{Nb}\frac{1}{1-b^{-1}}\sum_{a\in A_{1}}a^{-s_{0}-1}.$
We delay the discussion of the Riemann zeta function to the next section, and assume $s_{0}<1$ for now. There holds $\gamma_{1}\leq N(b-1)$, so we obtain as lower bound
$\lambda_{0}\log b>\sum_{a\in A_{1}}(1-s_{0}a^{-1})a^{-s_{0}}>0.$
This confirms that the residue is positive.
A simpler method is to start from the formula for $\lambda_{0}$ which follows from Equation (19) (recall the earlier observation that $K$ and $H$ have the same poles and residues for $\Re s=s_{0}$). We obtain:
$\lambda_{0}\log b=\sum_{a\in A_{1}}(a+1)^{-s_{0}}+\sum_{m=1}^{\infty}\frac{(s_{0})_{m}}{m!}N^{-1}b^{-m}\gamma_{m}^{\prime}H(s_{0}+m),$
which is visibly positive. More generally there holds, for $\lambda_{0,k}=\operatorname{Res}_{s_{0,k}}K(s)$:
$\lambda_{0,k}\log b=\sum_{a\in A_{1}}(a+1)^{-s_{0,k}}+\sum_{m=1}^{\infty}\frac{(s_{0,k})_{m}}{m!}N^{-1}b^{-m}\gamma_{m}^{\prime}H(s_{0,k}+m).$ (22)
We reverse-engineer the binomial series in the following manner, for any $s$ with $\Re s=\sigma>s_{0}-1$. Let us first consider this doubly infinite sum with non-negative terms:
$\sum_{m=1}^{\infty}\sum_{l=1}^{\infty}\sum_{a\in A}\frac{|(s)_{m}|}{m!}b^{-\sigma-m}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\frac{(b-1-a)^{m}}{(n+1)^{\sigma+m}}$
$\leq\frac{\Gamma(\sigma+1)}{|\Gamma(s+1)|}\sum_{m=1}^{\infty}\sum_{l=1}^{\infty}\sum_{a\in A}\frac{|s|}{|s+m|}\frac{(\sigma+1)_{m}}{m!}b^{-\sigma-m}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\frac{(b-1-a)^{m}}{(n+1)^{\sigma+m}}$
$\leq\frac{\Gamma(\sigma+1)}{|\Gamma(s+1)|}\sum_{l=1}^{\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\sum_{a\in A}(bn+b)^{-\sigma}\biggl(\Bigl(1-\frac{b-1-a}{bn+b}\Bigr)^{-\sigma-1}-1\biggr)$
$\leq\frac{\Gamma(\sigma+1)}{|\Gamma(s+1)|}\sum_{l=1}^{\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\sum_{a\in A}(bn+b)\bigl((bn+a+1)^{-\sigma-1}-(bn+b)^{-\sigma-1}\bigr).$
$\leq Nb^{-\sigma}\frac{\Gamma(\sigma+1)}{|\Gamma(s+1)|}\sum_{l=1}^{\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}(n+1)\bigl(n^{-\sigma-1}-(n+1)^{-\sigma-1}\bigr)$
$\leq Nb^{-\sigma}(\sigma+1)\frac{\Gamma(\sigma+1)}{|\Gamma(s+1)|}\sum_{l=1}^{\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}(n+1)n^{-\sigma-2}<\infty.$
Thus, it is finite. Hence the next computations are justified as soon as $\Re s>s_{0}-1$:
$\sum_{m=1}^{\infty}\frac{(s)_{m}}{m!}b^{-s}b^{-m}\gamma_{m}^{\prime}H(s+m)$
$=\sum_{m=1}^{\infty}\sum_{l=1}^{\infty}\sum_{a\in A}\frac{(s)_{m}}{m!}b^{-s-m}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\frac{(b-1-a)^{m}}{(n+1)^{s+m}}$
$=\sum_{l=1}^{\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\sum_{a\in A}(bn+b)^{-s}\left(\bigl(1-\frac{b-1-a}{bn+b}\bigr)^{-s}-1\right)$
$=\lim_{\ell\to\infty}\sum_{l=1}^{\ell}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\sum_{a\in A}\left((bn+a+1)^{-s}-(bn+b)^{-s}\right)$
Something special happens if $b^{s}=N$, i.e. if $s=s_{0,k}$, $k\in\mathbb{Z}$, as then $\sum_{a\in A}b^{-s}=Nb^{-s}=1$. So the computation gives:
$\sum_{m=1}^{\infty}\frac{(s)_{m}}{m!}N^{-1}b^{-m}\gamma_{m}^{\prime}H(s+m)$
$=\lim_{\ell\to\infty}\sum_{l=1}^{\ell}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=l}\left(\Bigl(\sum_{a\in A}(bn+a+1)^{-s_{0,k}}\Bigr)-(n+1)^{-s_{0,k}}\right)$
$=-\sum_{a\in A_{1}}(a+1)^{-s_{0,k}}+\lim_{\ell\to\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=\ell+1}(n+1)^{-s_{0,k}}.$
Combining with Equation (22) leads to the next proposition:
Proposition 5.1.
The residues $\lambda_{0,k}=\operatorname{Res}_{s_{0,k}}K(s)$ verify:
$\lambda_{0,k}=(\log b)^{-1}\lim_{\ell\to\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=\ell}\frac{1}{n^{s_{0,k}}}.$ (23)
Proof.
The chain of equations prior to the proposition, and leading to a telescopic series, proves from (22) that the residue of $H(s)$ at $s=s_{0,k}$ is:
$(\log b)^{-1}\lim_{\ell\to\infty}\mathop{\sum\nolimits^{\prime}}\limits_{\ell(n)=\ell}\frac{1}{(n+1)^{s_{0,k}}}.$
There only remains to explain why we can replace $n+1$ above with $n$. We use again the simple-minded upper bound $|n^{-s}-(n+1)^{-s}|\leq|s|n^{-\sigma-1}$, valid for $\sigma+1\geq 0$, in particular it is valid with $s=s_{0,k}$ for which $\sigma=s_{0}$. And as $\sum^{\prime}n^{-s_{0}-1}<\infty$, the contribution of the admissible integers of length $\ell$ goes to zero as $\ell$ goes to $\infty$. ∎
6. The case of the zeta function
Let us suppose that there are no restrictions on the digits, i.e. $N=b$. Then $s_{0}=1$ and $K(s)=\zeta(s)$. The formula of Proposition 5.1 says that
$\lambda_{0}=(\log b)^{-1}\lim_{\ell\to\infty}\sum_{b^{\ell-1}\leq n<b^{\ell}}n^{-1},$
which by a Riemann sum argument using $\int_{b^{-1}}^{1}\frac{\mathrm{d}t}{t}$ gives $\lambda_{0}=1$. More generally, at $s=1+k2\pi i/\log(b)$, $k\neq 0$, we need to evaluate
$\lim_{\ell\to\infty}\sum_{b^{\ell-1}\leq n<b^{\ell}}n^{-1-k2\pi i/\log(b)}$ $=\lim_{\ell\to\infty}b^{-\ell(1+k2\pi i/\log(b))}\sum_{b^{\ell-1}\leq n<b^{\ell}}(b^{-\ell}n)^{-1-k2\pi i/\log(b)}$
$=\int_{b^{-1}}^{1}t^{-1-k2\pi i/\log b}\mathrm{d}t=\left[\frac{t^{-k2\pi i/\log(b)}}{-k2\pi i/\log(b)}\right]_{b^{-1}}^{1}=0.$
This (fortunately…) confirms that there are no poles apart from $s=1$ on the line $\Re(s)=1$. Thus, using Equation (25), the other poles, if any, are at non-positive integers.
We continue the discussion later, but indulge here in a short digression about Equations (21) and (22). They give respectively:
$\log b$ $=\sum_{1\leq a<b}\frac{1}{a}+\sum_{m=1}^{\infty}(-1)^{m}b^{-m-1}(\sum_{1\leq a<b}a^{m})\zeta(m+1),$
$\log b$ $=\sum_{1\leq a<b}\frac{1}{a+1}+\sum_{m=1}^{\infty}\frac{\sum_{1\leq a<b}a^{m}}{b^{m+1}}(\zeta(m+1)-1).$
The second one has geometric convergence of ratio $(1-b^{-1})/2$, we can improve the quality of the first one this way:
$\log b$ $=\sum_{1\leq a<b}\frac{1}{a}+\sum_{m=1}^{\infty}\frac{(-1)^{m}a^{m}}{b^{m+1}}+\sum_{m=1}^{\infty}(-1)^{m}b^{-m-1}(\sum_{1\leq a<b}a^{m})(\zeta(m+1)-1)$
$=\sum_{1\leq a<b}\frac{1}{a}-\sum_{0\leq a<b}\frac{a}{b(a+b)}+\sum_{m=1}^{\infty}(-1)^{m}b^{-m-1}(\sum_{1\leq a<b}a^{m})(\zeta(m+1)-1)$
$=\sum_{1\leq a<2b}\frac{1}{a}-1+\sum_{m=1}^{\infty}(-1)^{m}\frac{\sum_{1\leq a<b}a^{m}}{b^{m+1}}(\zeta(m+1)-1).$
We can average the two representations of $\log b$ to have a sum involving only odd zeta values. Here is with $b=2$:
$\log 2$ $=\dfrac{5}{6}-\sum_{m=1}^{\infty}(-1)^{m-1}\dfrac{\zeta(m+1)-1}{2^{m+1}}$
$=\dfrac{1}{2}+\sum_{m=1}^{\infty}\dfrac{\zeta(m+1)-1}{2^{m+1}},$
$=\frac{2}{3}+\frac{1}{2}\sum_{p=1}^{\infty}\frac{\zeta(2p+1)-1}{4^{p}}.$
The last formula (average of the first two) is geometric with ratio $1/16$. A slower equivalent form is $\frac{2}{3}+\frac{1}{2}\sum_{n=2}^{\infty}(n(4n^{2}-1))^{-1}$.
Let us examine again (15) for the Riemann zeta function (and using $\ell=1$). So $N=b$ and the formula reads:
$\bigl(1-b^{1-s}\bigr)\zeta(s)=\sum_{1\leq a<b}a^{-s}+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}b^{-s-m}\gamma_{m}\zeta(s+m).$ (24)
Due to $(s)_{m}=s(s+1)\dots(s+m-1)$, the series extended from $m=2$ to $\infty$ (which is absolutely convergent for $\Re s>-1$) vanishes at $s=0$. And the term with $m=1$ gives a contribution equal to $b^{-1}\gamma_{1}=-(b-1)/2$ as we already know that the residue is $1$ at $s=1$. So, there is no pole but a finite value at $s=0$:
$(1-b)\zeta(0)=b-1-\frac{b-1}{2}=\frac{b-1}{2}\implies\zeta(0)=-\frac{1}{2}.$
Looking at Equation (24) for $\frac{3}{2}<\Re s<-\frac{1}{2}$, the terme with $m=1$ (evaluation at $s+1$) gives no pole, and $m=2$ does not either as the numerator vanishes at s=1s=-1. Proceeding inductively to $\frac{1}{2}-m<\Re s<\frac{1}{2}-m+1$ for $m>2$ to check if there is a pole at s=1ms=1-m, we again find that the $m$-th term of the series actually contributes no pole because the numerator vanishes. So the zeta function has only a single pole in the complex plane, located at $s=1$.
7. Residues (II)
We return to the general case.
Proposition 7.1.
Let $\lambda_{m,k}$ be the residue (perhaps vanishing) of $K(s)$ at $s_{m,k}=s_{0}-m+k2\pi i/\log b$. The following relations hold for $m\geq 1$:
$(1-b^{-m})\lambda_{m,k}=-N^{-1}\sum_{j=1}^{m}\frac{(m-s_{0,k})\dots(m-j+1-s_{0,k})}{j!}\frac{\gamma_{j}}{b^{j}}\lambda_{m-j,k}.$ (25)
If $\lambda_{0,k}=0$, i.e. if $s_{0,k}$ is not a pole, then $s_{m,k}$ is not a pole either, for every $m\geq 1$.
Proof.
The linear recurrence is a consequence of (15) (details left to reader; it is recommended to first replace the letter $m$ used as summation index in (15) by the letter $j$). And induction using (25) then shows that if $\lambda_{0,k}=0$ then $\lambda_{m,k}$ vanishes for every $m\geq 1$. ∎
The equation (25) is also valid for $m=0$ as the empty sum on the right-hand side has value zero.
We define new quantities $\mu_{m,k}$ such that the following relation holds:
$\lambda_{m,k}=\frac{-(-s_{0,k})_{m+1}}{m!}\mu_{m,k}=(-1)^{m}\frac{s_{m,k}(s_{m,k}+1)\dots(s_{m,k}+m)}{m!}\mu_{m,k}.$ (26)
But this definition fails for $k=0$ if $(-s_{0})_{m+1}$ can vanish, i.e. if $s_{0}$ is 0 (case $N=1$) or $1$ (case $N=b$). We will define later $\mu_{m,0}$ if $N=1$ and $N=b$ in an ad hoc manner, in order for the next Proposition to hold for them too. For $1<N<b$ one has in particular:
$\mu_{0,k}=\frac{\lambda_{0,k}}{s_{0,k}},\quad\mu_{1,k}=\frac{-\lambda_{1,k}}{s_{0,k}(s_{0,k}-1)}.$
Proposition 7.2.
The following recurrence formula holds, for $m\geq 1$:
$(1-b^{-m})\mu_{m,k}=-N^{-1}\sum_{j=1}^{m}\binom{m}{j}b^{-j}\gamma_{j}\mu_{m-j,k}\,.$ (27)
It can be written in the alternative form, with $\mu_{m,k}$ on both sides:
$\mu_{m,k}=\frac{b^{m}}{N}\sum_{j=0}^{m}\binom{m}{j}\frac{\gamma_{j}}{b^{j}}\mu_{m-j,k}\,.$ (28)
Remark 7.1.
For $N=1$, and $k=0$, equation (26) can not be used because $s_{0}=0$. For $k\neq 0$ one checks via a simple computation from Equation (11), which gives the value of $\lambda_{m,k}$, that
$\mu_{m,k}=\frac{(-1)^{m}}{2\pi ik}\qquad(N=1;k\neq 0).$
In order for Proposition 7.2 to hold also for $(\mu_{m,0})$, i.e. for the recurrence (27) to be verified, we can choose any constant $c$ and set $\mu_{m,0}=(-1)^{m}c$ because we know that this recurrence is verified by $((-1)^{m}(2\pi ik)^{-1})_{m\geq 0}$, for example with $k=1$. We choose $\mu_{m,0}=(-1)^{m}$.
For the Riemann zeta function, we define $(\mu_{m,0})$ to be the solution of (27) such that $\mu_{0,0}=1$. It turns out that then $\mu_{m,0}=B_{m}$, the $m$-th Bernoulli number (with $B_{1}=-\frac{1}{2}$). Indeed, the recurrence (28), in this $N=b$ case, is solved by the Bernoulli numbers due to the identity $(\mathsf{e}^{t}-1)^{-1}\sum_{a=0}^{b-1}\exp(b^{-1}at)=(\mathsf{e}^{b^{-1}t}-1)^{-1}$.
Proof of Proposition 7.2.
The details of the deduction of Equation (27) from Equations (25) and (26), if $1<N<b$, or if $N$ is arbitrary but $k$ is non-zero, are left to the reader. In the case $N=1$ and $N=b$ with $k=0$, the equation (27) is true due to the ad hoc definition of the sequence $(\mu_{m,0})$. The formula of (27) holds also with $m=0$, as an empty finite sum is defined to be zero. One uses $\gamma_{0}=N$ to relate the two ways of writing the recurrence. ∎
An immediate observation relative to (27) is that when we vary $k$ we keep the exact same recurrence relation. Hence:
Proposition 7.3.
For every $m\in\mathbb{N}$ and $k\in\mathbb{Z}$ there holds:
$\mu_{m,k}=\frac{\mu_{0,k}}{\mu_{0,0}}\mu_{m,0}.$ (29)
In other terms, once we know the residues on the horizontal axis, we don’t need to solve again (25) for $k\neq 0$, if we know $\lambda_{0,k}$.
Let us now consider the (formal, at this stage) normalized exponential generating function for the modified residues on the horizontal axis:
$B(t)=(\mu_{0,0})^{-1}\sum_{m=0}^{\infty}\frac{\mu_{m,0}}{m!}t^{m}.$
We compute the product:
$(\sum_{a\in A}\mathsf{e}^{b^{-1}at})B(t)$ $=\sum_{j=0}^{\infty}\frac{b^{-j}\gamma_{j}}{j!}t^{j}\sum_{q=0}^{\infty}\frac{\mu_{q,0}}{q!}t^{q}$
$=\sum_{m=0}^{\infty}(\sum_{j=0}^{m}\binom{m}{j}b^{-j}\gamma_{j}\mu_{m-j,0})\frac{t^{m}}{m!}$
$=\sum_{m=0}^{\infty}Nb^{-m}\mu_{m,0}\frac{t^{m}}{m!}=NB(b^{-1}t).$
This gives the functional equation:
$B(t)=\frac{N}{\sum_{a\in A}\mathsf{e}^{b^{-1}at}}B(b^{-1}t).$ (30)
We can now state the following somewhat striking corollary:
Theorem 7.4.
Let $E(t)$ be the moment generating function for the canonical probability measure with support the Cantor set of those real numbers in $[0,1]$ whose infinite $b$-radix expansion (or at least one of them, if there are two) uses admissible digits only. The exponential generating function $B(t)=(\mu_{0,0})^{-1}\sum_{m=0}^{\infty}\frac{\mu_{m,0}}{m!}t^{m}$ of the normalized residues on the real axis of the restricted Dirichlet series $K(s)$ is its multiplicative inverse:
$B(t)=\lim_{k\to\infty}\frac{N^{k}}{\sum\limits_{a_{1},\dots,a_{k}\in A}\mathsf{e}^{t\sum_{j=1}^{k}b^{-j}a_{j}}}=\Bigl(\prod_{j=1}^{\infty}\frac{\sum_{a\in A}\mathsf{e}^{b^{-j}at}}{N}\Bigr)^{-1}\;.$ (31)
It is a meromorphic function in the entire complex plane, as inverse of an entire function. It verifies $\Re t<0\implies|B(t)|>1$.
Remark 7.2.
The reference to poles on the real axis is a bit vague, as it allows for a vanishing residue. The precise formulation is given by the formula defining $B(t)$ and the prior definitions. When we refer to $b$-radix expansion, we mean an infinite one: so if 0 is not an admissible digit, and b1b-1 isn’t either, “$b$-decimal” numbers are not in the alluded-to Cantor set even if their finite expansion uses only digits from $A$. And of course, in the case of the Riemann zeta function, the “Cantor” set is the full unit interval $[0,1]$.
Remark 7.3.
In the case $b=N$, which corresponds to the Riemann zeta function $\zeta(s)$, the function $E(t)$ is $\int_{0}^{1}\mathsf{e}^{tx}\mathrm{d}x=\frac{\mathsf{e}^{t}-1}{t}$. Hence $B(t)=\frac{t}{e^{t}-1}$ is indeed the exponential generating function of the $\mu_{m,0}/\mu_{0,0}$ quantities in that case, as we have argued earlier that $\mu_{m,0}=B_{m}$ due to their ad hoc definition as the solution of Equation (28) with $\mu_{0,0}=1$. This makes, in general, the ratios $\mu_{m,0}/\mu_{0,0}$ (which are rational numbers) some sort of generalized Bernoulli numbers associated with $b>1$ and $\subset[\![0,b-1]\!], AA\neq\emptyset$, $A\neq\{0\}$.
Proof of Theorem 7.4.
We first show that the infinite product
$E(t)=\prod_{j=1}^{\infty}\frac{\sum_{a\in A}\mathsf{e}^{b^{-j}at}}{N}$
converges everywhere in the complex plane and defines an entire function. For $N=1$ and $f>0$ the allowed digit, this clearly is the case with $E(t)=\exp(\frac{f}{b-1}t)$. If $N>1$, it is also clearly the case from the fact that
$\frac{\sum_{a\in A}\mathsf{e}^{b^{-j}at}}{N}=1+O_{t\to 0}(t).$
And, for any bounded region, only finitely many terms may have zeros in it, and removing those terms we get a uniformly convergent non-zero product. Zeros of the entire function $E(t)$ are of the shape $b^{j}z$, $j\geq 1$, where $z$ is in finitely many additive cosets of $2\pi i\mathbb{Z}$, corresponding to $\mathsf{e}^{z}$ being a non-zero root of the polynomial $P_{A}(z)=\sum_{a\in A}z^{a}$. The function $E$ is of finite exponential type f/(b1)f/(b-1) but we don’t need this. For $\Re t<0$, each term of the product is of complex modulus less than $1$, so $|E(t)|<1$.
Let $r>0$ be the smallest modulus of a root of $E$ (or $r=\infty$ for $N=1$), and consider the power series expansion of $B(t)=E(t)^{-1}$:
$B(t)=1+\sum_{j=1}^{\infty}\frac{c_{j}}{j!}t^{j},$
which exists and is convergent for $|t|<r$. We set $c_{0}=1$. By construction of $E(t)$ there holds
$B(t)=\frac{N}{\sum_{a\in A}\mathsf{e}^{b^{-1}at}}B(b^{-1}t).$
Reversing the steps of an earlier calculation this means that the sequence $(c_{m})_{m\geq 0}$ shares the same equivalent recurrence formulas (27) and (28) as does the sequence $(\mu_{m,0}/\mu_{0,0})_{m\geq 0}$. As they have the same starting point at $m=0$, this establishes that $c_{m}=\mu_{m,0}/\mu_{0,0}$.
To complete the proof we explain briefly the connection with Cantor sets. Consider infinitely many identically distributed and independent random variables $X_{j}$, $j\geq 1$, each probability $1/N$ to take any given value $a\in A$. It is well-known that, if $\Omega$ is the probability space, and $x:\Omega\to[0,1]$ is the map
$x(\omega)=\sum_{j=1}^{\infty}\frac{X_{j}(\omega)}{b^{j}},$
(which is maybe not one-to-one if $A$ contains both 0 and b1b-1, but this is benign) then the push-forward under $x$ of the probability measure to $[0,1]$ gives a Borel measure, which is Lebesgue measure if $N=1$, else is singular to Lebesgue measure and supported by the Cantor set whose first construction step is to replace $[0,1]$ by the union of the intervals $b^{-1}a,b^{-1}(a+1)]$, $a\in A$, then repeat the operation in each of them, etc…. For any complex $t$, the expectation value of the random variable $\mathsf{e}^{tx}$ on $\Omega$ becomes the integral of the function $x\mapsto\mathsf{e}^{tx}$ against that push-forward probability measure. If we compute it on $\Omega$ as the limit from approximation by simple functions constant on cylinder sets defined by finitely many conditions $X_{1}=a_{1}$, …, $X_{k}=a_{k}$, we obtain that this expansion is the limit as $k$ goes to infinity of
$N^{-k}\sum\limits_{a_{1},\dots,a_{k}\in A}\mathsf{e}^{t\sum_{j=1}^{k}b^{-j}a_{k}}=\prod_{j=1}^{k}\frac{\sum_{a\in A}\mathsf{e}^{b^{-j}at}}{N}.$
In other terms it is exactly $E(t)$. ∎
8. Values at non-positive integers
In this final section we study the values of $K(s)$ at the non-positive integers. We shall assume $N>1$, as the non-positive integers are poles if $N=1$. We also assume $N<b$, as the pole at $s=1$ of the Riemann zeta function modifies the details of the discussion.
Consider Equation (15) and substitute $s=0$. All contributions of the series on the right-hand side of this equation vanish. It follows:
$K(0)=\frac{-N_{1}}{N-1}.$
We will discover that it is important whether $N_{1}=N$ or $N_{1}=N-1$.
Proposition 8.1.
The sequence $(K(-m))_{m\geq 0}$, assuming $1<N<b$, consists of rational numbers verifying $K(0)=-N_{1}/(N-1)$ and the following recurrence:
$m\geq 1\implies(b^{-m}-N)K(-m)=\frac{\gamma_{m}}{b^{m}}+\sum_{j=1}^{m}\binom{m}{j}\frac{\gamma_{j}}{b^{j}}K(-m+j).$ (32)
If 0 is an admissible digit, i.e. if K(0)=1K(0)=-1, then K(m)=0K(-m)=0 for every $m\geq 1$. If 0 is not an admissible digit, i.e. if $N_{1}=N$, equivalently if K(0)<1K(0)<-1, the exponential generating function $C(t)=\sum_{m=0}^{\infty}K(-m)\frac{t^{m}}{m!}$ verifies the following functional equation:
$C(b^{-1}t)=\alpha_{A}(b^{-1}t)+\alpha_{A}(b^{-1}t)C(t),$ (33)
where $\alpha_{A}(t)=\sum_{a\in A}\mathsf{e}^{at}$. It has a positive radius of convergence and is a meromorphic function in the entire complex plane. It is given “explicitly” as
$C(t)=\sum_{m=0}^{\infty}K(-m)\frac{t^{m}}{m!}=-1-\sum_{k=1}^{\infty}\frac{1}{\prod_{j=1}^{k}\alpha_{A}(b^{-j}t)}\,.$ (34)
At every location except at the elements of $\{b^{j}z,j\geq 1,\alpha_{A}(z)=0\}$, the series at the right-hand side of (34) converges geometrically.
Proof.
Equation (32) is a corollary to (15) (details left to reader; it is recommended though to replace letter $m$ in (15) by letter $j$ at the very start). As $K(0)$ is rational, all the K(m)K(-m)’s are, too. The equation determining K(1)K(-1) is
$(b^{-1}-N)K(-1)=b^{-1}\gamma_{1}+b^{-1}\gamma_{1}K(0).$
So, if K(0)=1K(0)=-1, we get K(1)=0K(-1)=0. But then
$(b^{-2}-N)K(-2)=b^{-2}\gamma_{2}+b^{-2}\gamma_{2}K(0)=0.$
And by induction K(m)=0K(-m)=0 for every $m\geq 1$.
We now suppose that $N_{1}\neq N-1$, i.e. $0\notin A$ and $N_{1}=N$. We rewrite (32) in the equivalent form (as $\gamma_{0}=N$):
$b^{-m}K(-m)=\frac{\gamma_{m}}{b^{m}}+\sum_{j=0}^{m}\binom{m}{j}\frac{\gamma_{j}}{b^{j}}K(-m+j).$
Observe that this equation holds true also for $m=0$. This is due to $\gamma_{0}=N$ and $N=N_{1}$. So, the formal power series $C(t)$ verifies:
$C(b^{-1}t)=\sum_{m=0}^{\infty}\frac{\gamma_{m}}{b^{m}}\frac{t^{m}}{m!}+C(t)\sum_{m=0}^{\infty}\frac{\gamma_{m}}{b^{m}}\frac{t^{m}}{m!}.$
As $\sum_{m=0}^{\infty}\gamma_{m}\frac{t^{m}}{m!}=\alpha_{A}(t)$ we recognize the functional equation (33). Conversely, this identity of power series leads back to the recurrence (this again crucially uses $N=N_{1}$; we can write a functional equation in the general case, but as its solution if $N_{1}<N$ is already known we skip that).
Let us now directly examine the expression given in the right-hand side of (34). The denominators vanish at the points of the set $\{b^{j}z,j\geq 1,\alpha_{A}(z)=0\}$ whose intersections with bounded regions are finite. Take any large closed disk and consider the right-hand side of (34) but removing from it all occurrences of those finitely many functions $\alpha_{A}(b^{-1}t)$, …, $\alpha_{A}(b^{-k}t),\dots$ having zeros in this disk. The remaining expression is then everywhere defined, and as $\alpha_{A}(b^{-j}t)$ converges uniformly on such bounded region toward $N>1$ as $j\to\infty$, the series converges geometrically, and defines an analytic function. Restoring the removed factors we obtain that the expression converges everywhere away from the poles and defines a meromorphic function $D(t)$. Let $r>0$ be the smallest modulus of a root of $\alpha_{A}(b^{-1}t)$, we can expand $D(t)$ for $|t|<r$ into a convergent power series $\sum_{m=0}^{\infty}d_{m}t^{m}/m!$. The function $D(t)$ verifies the functional equation (33). Note in passing that this is not an homogeneous equation and the value $D(0)$ is constrained by it, but we can also compute it from the series as $\sum_{k=0}^{\infty}N^{-k}=-(1-1/N)^{-1}=-N/(N-1)=K(0)$. The coefficients $(d_{m})_{m\geq 0}$ verify the same recurrence and are thus identical with the $(K(-m))_{m\geq 0}$. This completes the proof. ∎